Journalist Report – February 5th

Crew 274 Journalist Report 2-5-2023

Sol: 7

Author: Tony DiBernardo, Crew Journalist

Title: The Long Hike

Crew 274 woke up this morning at 7:30am to the song “Take Me Home, Country Roads.” by John Denver. The crew got their surveys done, then made sure to get a decent breakfast before their 3-hour EVA.

EVA Activity

Today, HSO Pender and Hab Specialist O’Hara conducted a 3-hour EVA past the supply cache and back. The original EVA plan was to hike an hour to the supply cache, resupply their personal food and water from the cache, then hike 30 more minutes before turning around and hiking 1.5 hours back to the hab. They reached the cache early at the 47-minute mark due to them taking flat roads rather than the originally planned rocky terrain, but eventually found a more consistent and realistic pace on the way back. Supporting their EVA was Commander Guthrie and Journalist DiBernardo. They performed a “leap-frog” maneuver during the EVA, where they would drive the rover past the hiking crew and stop just before losing sight of them. Then, Pender and O’Hara would hike past the parked rover and just before they got out of site, Guthrie and DiBernardo would catch up in the rover and drive past them, only stopping just before losing sight again. This technique was used to reserve rover battery power as well as ensure a safe EVA for the crew on their longest hike of the mission. The EVA was a success, and apart from a couple of cars and hikers driving and passing by, the EVA went as expected.


Following the exciting germination of one LDEF seed yesterday, Greenhab Officer Hines, did his daily collection and observations of germination from the other seeds.. While no additional progress has been noted in other LDEF seeds, multiple sets of seeds that are yet to germinate were removed from the growth tent, placed in wet paper towels, and moved them into the GreenHab to utilize the higher temperatures and humidity in order to potentially germinate leading into the second half of the mission. Additionally, the cress microgreens have continued to make significant progress, with the red cabbage and broccoli making gradual yet minimal continued germination.


Crew astronomers Pena and Loy’s observations of heliophysics phenomena were significantly impaired by the site’s cloudy weather. Nevertheless, some useful data on solar dark spots and solar prominences were collected and analyzed. These observations will be very useful for the astronomer’s analysis of peak solar cycle 25. So far, 220,000 image points have been collected, and 50,000 images have been processed and analyzed.


Cloudy weather conditions canceled all planned observations of HADS Variable Star V0799 AUR. Crew Astronomers instead continued their analysis of their current image inventory, stacking and analyzing 12 separate 60-second images in the visual filter. So far, HADS Variable Star V0799 AUR is showing signs of dimming, which provides good data to the astronomers’ variable star report.

Stress Mitigation Study

Last night, the crew began their participation in Crew Engineer Lojeks’ study of tracking the stress levels of six of the crew members while practicing breathing and mindfulness techniques as a potential mitigation of stress in analog environments. He spoke with the crew about why he is pursuing this study, how to enter into this type of meditation and guided them through a 15-minute breathing exercise.

Copyright © The Mars Society. All rights reserved. | Main Site